OpenStudy (anonymous):

find all the solutions to x^4 - 16 = 0...I got -2, 2...correct?

OpenStudy (dan815):

(x^2 + 4) * (x^2 - 4) x^2 = sqrt-4 = +/- 2i if u looking for complex solutions too

OpenStudy (dan815):

(x^2 + 4) * (x^2 - 4) sqrt x^2 = sqrt-4 = +/- 2i if u looking for complex solutions too